• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack (Techs)

    Log entry time 17:03:45 on November 13, 2008

    Entry number 248653

    keyword=Right spectrometer status

    The Right Spectometer is under vacuum. Right Q1 is not usable, the ZFCT readout has been sent out for repair. Right Dipole has been turned on, but not ramped up to an appreciable current. Right Q2 is good to go. Right Q3 had about 60VAC from the output flags to ground, and from C1 to ground. One of the AC power wires for the lead flag heater resistors was pinched under the head of a screw on the flag. The wire was freed and some poly tubing fit around it for a little protection. The 60VAC is no longer on the flag at the power supply and the heater still seems to be working. The supply is still locked out till final testing is done.